# Microquasar jet-cocoon systems as PeVatrons

### v1

@@ Abstract
The origin of Galactic cosmic rays (CRs), particularly around the knee region ($\sim$3 PeV), remains a major unsolved question. Recent observations by LHAAASO suggest that the knee is shaped mainly by protons, with a transition to heavier elements at higher energies. Microquasars – compact jet-emitting sources – have emerged as possible PeV CR accelerators, especially after detections of ultrahigh-energy gamma rays from these systems. We propose that the observed proton spectrum (hard below a few PeV, steep beyond) arises from the reacceleration of sub-TeV Galactic CRs via shear acceleration in large-scale microquasar jet-cocoon structures. Our model also naturally explains the observed spectrum of energies around a few tens of PeV by summing up heavier nuclei contributions. Additionally, similar reacceleration processes in radio galaxies can contribute to ultrahigh-energy CRs, bridging Galactic and extragalactic origins. Combined with low-energy CRs from supernova remnants and galaxy clusters around the second knee region, this scenario could provide a unified explanation for CRs across the entire energy spectrum.
@@ I Introduction
The origin of cosmic rays (CRs), mainly composed of protons, heliums, and heavier nuclei, is a long-standing puzzle in astrophysics Antoni et al. (2005); Kotera and Olinto (2011); Kachelriess and Semikoz (2019). The energy spectrum of CRs spans over a wide range from $10^{9}\rm~{}eV$ to $10^{20}\rm~{}eV$, roughly following a power-law spectrum $\Phi_{\rm CR}\propto E^{-\gamma}$. The spectrum exhibits a distinctive feature with a change in the power law spectral index from $\gamma\sim 2.7$ to $\gamma\sim 3$ at the energy around $4\rm~{}PeV$, known as knee Antoni et al. (2005); Amenomori et al. (2008); Cao et al. (2024a), discovered by Kulikov and Khristiansen in 1958 Kulikov and Khristiansen . At energies around $\sim 100\rm~{}PeV$, the spectral index changed from $\gamma\sim 3$ to $\gamma\sim 3.3$, called the “second knee” (e.g., Nagano et al., 1992; Abbasi et al., 2018). The spectral index changes back to $\gamma\sim 2.5$ at the ankle around 4 EeV (e.g., Bird et al., 1993), and the flux is largely suppressed at a few tens of EeV Abbasi et al. (2008); Abraham et al. (2008). The knee is believed to be formed due to the limited maximum rigidity of Galactic CR accelerators Hillas (2005), or reflecting a spectral break at the same rigidity Cardillo et al. (2015). The knee could also be formed due to the change of the energy dependent diffusion coefficient in the propagation of CRs Giacinti et al. (2015), while CRs between the knee and ankle should be marked as the transition region from Galactic to extragalactic CRs Gaisser (2006). CRs with energy above EeV are regarded as ultrahigh-energy cosmic rays (UHECRs), which have an extragalactic origin as supported by the observation of the large-scale anisotropy, where the dipole direction points in the opposite direction of the Galactic center Aab et al. (2018).
@@
CR composition is the key to identifying the origins. The general pattern of CR abundances around a few GeV per nucleon resembles that of the interstellar medium or primordial solar composition, with peaks in the CNO and Fe region Drury (2018). Recently, a knee-like feature around $3\rm~{}PeV$ is also revealed on the measured all-particle spectrum of CRs. The logarithm of the mean mass of CRs measured by the Large High Altitude Air Shower Observatory (LHAASO) showed trends toward a heavier composition below several hundred TeV and above several PeV, while it is a light composition dominated around the knee, and the turning point is consistent with the position of the knee at energy around $3\rm~{}PeV$ Cao et al. (2024a). In particular, LHAASO reported the first high-purity proton spectrum, implicating the knee at the energy around $3\rm~{}PeV$ is related to the change of the proton spectrum Cao et al. (2025). The observed hardening feature above several hundreds of TeV implicates the emergence of a new component that has the maximum energy at least up to a few PeV Cao et al. (2025).
@@
The physical origin of the knee is still unknown and should be related to Galactic PeVtrons. SNRs are believed to be responsible for CRs at least up to the knee Reynolds (2008); Blasi (2013). However, SNRs as PeVatrons are challenged by both theoretical and observational studies Bell (2004, 2005); Gabici et al. (2019); Giuliani and Cardillo (2024), potentially implying existence of another type of PeVatrons. Potential candidates for PeVtrons include massive stars and stellar clusters Casse and Paul (1982); Cesarsky and Montmerle (1983); Aharonian et al. (2019), pulsars and pulsar winds Bednarek and Protheroe (2002); Giller and Lipski (2002); Bednarek and Bartosik (2004); Ohira et al. (2018), superbubbles Bykov et al. (2020), hypernovae Sveshnikova, L. G. (2003) and interacting supernovae Murase et al. (2014); Inoue et al. (2021), and isolated black holes Ioka et al. (2017); Kimura et al. (2025). The detection of very high-energy (VHE) gamma rays has shown that microquasars are promising sites for accelerating CRs around PeV and above, indicating potential sources of Galactic PeV CRs Cao et al. (2024b). Microquasars are X-ray binaries with relativistic jets Mirabel and Rodriguez (1994) and display phenomenology similar to active galactic nuclei (AGNs), although the size of the system is orders of magnitude different. It is suggested that shear and one-shot acceleration in the large-scale jet of radio galaxies can be responsible for the origin of UHECRs Kimura et al. (2018a); Caprioli (2015); Mbarek and Caprioli (2019); Mbarek et al. (2023, 2025). This motivates us to investigate whether the shear acceleration by microquasar jets can be Galactic PeVatrons. Previous studies also suggest microquasars as PeVatrons Cooper et al. (2020); Kimura et al. (2021); Yue et al. (2024), but these studies consider CR acceleration by shocks in jets or magnetic reconnections in accretion flows.
@@
The paper is organized as follows. In Section II, we study the physics of Galactic microquasars and the shear acceleration process in the jet-cocoon system of microquasars. In Section III, we propagate the escaped CRs from microquasars and compare the results with observations. In Section IV, we combine CRs from Galactic SNRs and contributions from the shear acceleration process from extragalactic clusters of galaxies and radio galaxies, we proposed an overall picture of the origin of CRs on a wide energy range from GeV to $\sim 100\rm~{}EeV$. We give a summary and discussion in Sec. V.
@@ II.1 The physics of microquasar jets
Recently, Ref. Cao et al. (2024b) reported detection of gamma rays above 100 TeV from five microquasars, three of which (SS 433, V4641 Sgr, and GRS 1915+105) exhibit extended gamma-ray morphology whose emission zone is as large as $\sim 50-100$ pc . These gamma-rays are most likely originating from the radiation of nonthermal particles accelerated in the extended jets Alfaro et al. (2017); Aharonian et al. (2024); Cao et al. (2024b). The other two (Cyg X-1 and MAXI J1820+070) are consistent with point source, possibly suggesting a compact emission region Fang et al. (2024); Kuze et al. (2025), although the best fit positions of gamma-ray sources have slight offsets from those of X-ray binaries.
SS 433 is a binary system consisting of a compact object and a supergiant star, with a pair of jets moving in opposite directions with launch velocity $v\sim 0.26\rm~{}c$, where $c$ is the speed of light Abell and Margon (1979). The extended two-sided jets with $\sim 40$ pc are detected in X-rays Safi-Harb and Ögelman (1997), and radio lobes surrounding the system extends up to $\sim 150$ pc Dubner et al. (1998); Sakemi et al. (2023). VHE gamma-ray emission from SS 433 has been detected at the jet lobes, which are spatially coincident with the X-ray hotspots Abeysekara et al. (2018); Aharonian et al. (2024); Cao et al. (2024b); Alfaro et al. (2024a). The observed gamma-ray emission is consistent with inverse Compton emission from a single population of electrons, and the magnetic field strength inside the jet is $\sim 20\rm~{}\mu G$. Refs. Sudoh et al. (2020); Kimura et al. (2020) pointed out that nonthermal electrons accelerated at SS 433 jets achieves PeV energies with a small Bohm factor, based on multiwavelength modeling. In addition, Ref. Ohmura et al. (2021) performed magnetohydrodynamics (MHD) simulations of the propagation of the jet in the surrounding medium, where the jet is spheroidal at early times and the wings are developed by the broadening spherical cocoon.
V4641 Sgr is another interesting microquasar who emits gamma-rays above 100 TeV. The morophorogy of the gamma-ray signal is clear two-sided lobe structure, and the size of the lobe is roughly $\sim 100$ pc Cao et al. (2024b); Alfaro et al. (2024b). Spatially extended X-ray signature is also detected from this objects, although the X-ray signal is concentrated on the central region within $\sim 15$ pc from the central black hole (Suzuki et al., 2025). The multiwavelength modeling of this system indicates $B\gtrsim 10\rm\mu G$ at the X-ray emission region.
GRS 1915+105 is a highly variable object that exhibits super-luminal motion with extended jets of sub-pc scales Fender et al. (1999). Its Eddington ratio used to be $\sim 0.1-1$ (Done et al., 2004) and suddenly decreased to 0.01 in 2018 Koljonen and Hovatta (2021). The gamma-ray source detected by LHAASO exhibits an extension of $\sim 50$ pc Cao et al. (2024b), which is consistent with the two-sided lobe structure detected by ALMA Tetarenko et al. (2018).
@@
×$\times{\left(\frac{L_{j}}{{10}^{39}~{}{\rm erg/s}}\right)}^{1/2}\left(\frac{\Gamma}{1.05}\right)^{-1}\left(\frac{\beta}{0.3}\right)^{1/2}$ (1)
where $Z$ is the nuclear charge, $\epsilon_{B}$ is the energy fraction carried by the magnetic field, $L_{j}$ is the total isotropic-equivalent luminosity of the jet, $\Gamma$ is the Lorentz factor, and $\beta$ is the characteristic velocity normalized by the speed of light. In our Galaxy, there are likely a few dozens of BH X-ray binaries with an X-ray luminosity of $L_{X}\gtrsim 10^{38}\rm~{}erg~{}s^{-1}$, each exhibiting sufficient luminosity, $L_{j}\sim 10^{39}\rm~{}erg~{}s^{-1}$, to potentially accelerate CRs up to the knee energy range. Here, we use frequently used assumptions of $L_{X}\sim 0.1\dot{M}c^{2}$ and $L_{j}\sim\dot{M}c^{2}$ to convert $L_{X}$ to $L_{j}$, where $\dot{M}$ is the mass accretion rate onto the BH.
@@ II.2 Shear acceleration in microquasar jets
Shear acceleration, a class of the Fermi acceleration, has been suggested to be an efficient mechanism to accelerate CRs, where charged particles gain energy inside an ordered shear velocity field Berezhko and Krymskii (1981); Earl et al. (1988); Ostrowski (1990); Webb (1989); Rieger (2019). It has been shown that the discrete shear acceleration Ostrowski (1990) occurring on the surface of the jet-cocoon system in radio galaxies may accelerate low-energy “seed” galactic CRs to the UHE energy range, and the predicted spectrum and composition are consistent with observations Kimura et al. (2018a); Seo et al. (2024).
In AGNs, the reacceleration may occur through the “one-shot” acceleration, where the “seed” CRs receive a boost of a factor of $\Gamma^{2}$ Gallant and Achterberg (1999), and it is possible for powerful radio galaxies with a Lorentz factor of $\Gamma\gtrsim 30$ to accelerate “seed” galactic CRs with energy $E\lesssim 10^{17}\rm~{}eV$ to the UHE energy range Caprioli (2015); Mbarek and Caprioli (2019).
We consider the shear acceleration by cylindrical jets and cocoon as discussed in Kimura, Murase and Zhang (hereafter KMZ18) Kimura et al. (2018a). Based on discussion in Section II.1, we assume that the radius of the jet is $r_{\rm jet}=3\rm~{}pc$, the length is $l_{\rm jet}=100\rm~{}pc$, the jet velocity normalized by the speed of light is $\beta_{j}=0.3$ (corresponding to Lorentz factor $\Gamma_{j}\simeq 1.05$), the radius of the cocoon is $r_{\rm coc}=60\rm~{}pc$. The coherence length inside the cocoon is set to $l_{\rm coh}\sim 0.03r_{\rm coc}\sim 1.8\rm~{}pc$. The mean free path of CRs is given by $\lambda\sim l_{\rm coh}(E/E_{\rm coh})^{\delta}$, where $\delta=1/3$ for Kolmogorov turbulence and $E_{\rm coh}=ZeB_{\rm coc}l_{\rm coh}\sim 17Z(B_{\rm coc}/10{\rm~{}\mu G})(l_{\rm coh}/1.8\rm~{}pc)\rm~{}PeV$. CRs will undergo a discrete shear acceleration when the CR mean free path is longer than the width of the sheath layer, $\lambda\gtrsim r_{\rm sl}$. For $r_{\rm sl}\sim f_{\rm sl}r_{\rm jet}\sim 0.06(f_{\rm sl}/0.02)\rm~{}pc$, the injection energy is given by $\lambda\sim r_{\rm sl}$, which can be rewritten as
@@
$\sim$ $0.6\left(\frac{B_{\rm coc}}{\rm 10\mu G}\right)\left(\frac{r_{\rm sl}}{0.06\rm~{}pc}\right)^{3}\left(\frac{l_{\rm coh}}{\rm 1.8pc}\right)^{-2}\rm~{}TeV.$
Thus, CRs with energies above several hundreds of GeV can undergo a discrete shear acceleration mechanism. We should note that the width of the shear layer and injection energy to the discrete shear acceleration are still uncertain, and $f_{\rm sl}\sim 1$ is suggested by recent high-resolution MHD simulations when the kink instability is fully developed Wang et al. (2024).
Based on KMZ18 Kimura et al. (2018a), by balancing the acceleration and confinement timescales, the peak energy in the energy spectrum of escaping CR protons can be estimated to be
@@
$\simeq$ $4.3{\rm~{}PeV}\left(\frac{B_{\rm coc}}{10\mu\rm G}\right)\left(\frac{l_{\rm coh}}{1.8\rm~{}pc}\right)^{1/2}\left(\frac{r_{\rm jet}}{3\rm pc}\right)^{1/2}\left(\frac{\beta_{\rm jet}}{0.3}\right)\zeta,$
@@
×$\times\left(\frac{r_{\rm coc}}{60\rm~{}pc}\right)^{-1}\left(\frac{v_{\rm exp}}{3000\rm~{}km~{}s^{-1}}\right),$
@@
×$\times\left(\frac{r_{\rm coc}}{60\rm~{}pc}\right)^{-1}\left(\frac{B_{\rm coc}}{10\mu\rm G}\right)\left(\frac{l_{\rm coh}}{1.8\rm~{}pc}\right)^{1/2}\left(\frac{\beta_{\rm jet}}{0.3}\right)\zeta\rm~{}erg~{}s^{-1}.$
@@
Table 1: Model parameters used in Monte Carlo simulations.
We perform Monte Carlo simulations with the code used in KMZ18 Kimura et al. (2018a) with parameters given in Table 1, and the resulting energy spectra for escaping CR protons are shown in Fig. 2. As seen in the figure, our shear acceleration scenario can achieve spectral peak around PeV energies with reasonable parameters (see Section V for discussion of parameter choices). The escaping CR proton spectrum show a hard spectrum below the peak and a steep power-law decay beyond the peak.
Figure 2: The energy spectrum of escaping CRs predicted in the shear acceleration mechanism from Monte Carlo simulations. The parameters are listed in Table 1. The above plot is multiplied with $\rm exp(-\mathcal{R}/\mathcal{R}_{\rm max})$, where $\mathcal{R}_{\rm max}=4\times 10^{7}\rm~{}GeV$ is the maximum acceleration energy from microquasar jets.

### v2

@@ Abstract
The origin of Galactic cosmic rays (CRs), particularly around the knee region ($\sim$3 PeV), remains a major unanswered question. Recent observations by LHAASO suggest that the knee is shaped mainly by protons, with a transition to heavier elements at higher energies. Microquasars – compact jet-emitting sources – have emerged as possible PeV CR accelerators, especially after detections of ultrahigh-energy gamma rays from these systems. We propose that the observed proton spectrum (hard below a few PeV, steep beyond) arises from the reacceleration of sub-TeV Galactic CRs via shear acceleration in large-scale microquasar jet-cocoon structures. Our model also naturally explains the observed spectrum of energies around a few tens of PeV by summing up heavier nuclei contributions. Additionally, similar reacceleration processes in radio galaxies can contribute to ultrahigh-energy CRs, bridging Galactic and extragalactic origins. Combined with low-energy CRs from supernova remnants and galaxy clusters around the second knee region, this scenario could provide a unified explanation for CRs across the entire energy spectrum.
@@ I Introduction
The origin of cosmic rays (CRs), mainly composed of protons, heliums, and heavier nuclei, is a long-standing puzzle in astrophysics Antoni et al. (2005); Kotera and Olinto (2011); Kachelriess and Semikoz (2019). The energy spectrum of CRs spans over a wide range from $10^{9}$ to $10^{20}\rm~eV$, roughly following a power-law spectrum $\Phi_{\rm CR}\propto E^{-\gamma}$. The spectrum exhibits a distinctive feature with a change in the power-law spectral index from $\gamma\sim 2.7$ to $\gamma\sim 3$ at the energy around $4\rm~PeV$, known as the kneeknee Antoni et al. (2005); Amenomori et al. (2008); Cao et al. (2024a), discovered by Kulikov and Khristiansen in 1958 Kulikov and Khristiansen (1958). At energies $\sim 100\rm~PeV$, the spectral index changed from $\gamma\sim 3$ to $\gamma\sim 3.3$, called the “second knee” (e.g., Nagano et al., 1992; Abbasi et al., 2018). The spectral index changes back to $\gamma\sim 2.5$ at the ankle around 4 EeV (e.g., Bird et al., 1993), and the flux is largely suppressed at a few tens of EeV Abbasi et al. (2008); Abraham et al. (2008). The knee is believed to be formed due to the limited maximum rigidity of Galactic CR accelerators Hillas (2005), or reflecting a spectral break at the same rigidity Cardillo et al. (2015). The knee could also be formed due to the change of the energy-dependent diffusion coefficient in the propagation of CRs Giacinti et al. (2015), while CRs between the knee and ankle should be marked as the transition region from Galactic to extragalactic CRs Gaisser (2006). CRs with energy above EeV are regarded as ultrahigh-energy cosmic rays (UHECRs), which have an extragalactic origin as supported by the observation of the large-scale anisotropy, where the dipole direction points in the opposite direction of the Galactic Center Aab et al. (2018).
@@
CR composition is the key to identifying the origins. The general pattern of CR abundances around a few GeV per nucleon resembles that of the interstellar medium or primordial solar composition, with peaks in the carbon-nitrogen-oxygen (CNO) and Fe region Drury (2018). Recently, a kneelike feature around $3\rm~PeV$ was also revealed on the measured all-particle spectrum of CRs. The logarithm of the mean mass of CRs measured by the Large High Altitude Air Shower Observatory (LHAASO) showed trends toward a heavier composition below several hundred TeV and above several PeV, while it is a light composition dominated around the knee, and the turning point is consistent with the position of the knee at energy around $3\rm~PeV$ Cao et al. (2024a). In particular, LHAASO reported the first high-purity proton spectrum, implicating the knee at energy around $3\rm~PeV$ is related to the change of the proton spectrum Cao et al. (2025). The observed hardening feature above several hundreds of TeV implicates the emergence of a new component that has maximum energy at least up to a few PeV Cao et al. (2025).
@@
The physical origin of the knee is still unknown and should be related to Galactic PeVatrons. Supernova remnants (SNRs) are believed to be responsible for CRs at least up to the knee Reynolds (2008); Blasi (2013). However, SNRs as PeVatrons are challenged by both theoretical and observational studies Bell (2004, 2005); Gabici et al. (2019); Giuliani and Cardillo (2024), potentially implying the existence of another type of PeVatrons. Potential candidates for PeVatron include massive stars and stellar clusters Casse and Paul (1982); Cesarsky and Montmerle (1983); Aharonian et al. (2019), pulsars and pulsar winds Bednarek and Protheroe (2002); Giller and Lipski (2002); Bednarek and Bartosik (2004); Ohira et al. (2018), superbubbles Bykov et al. (2020), hypernovae Sveshnikova, L. G. (2003) and interacting supernovae Murase et al. (2014); Inoue et al. (2021), and isolated black holes Ioka et al. (2017); Kimura et al. (2025). The detection of very-high-energy (VHE) gamma rays has shown that microquasars are promising sites for accelerating CRs around PeV and above, indicating potential sources of Galactic PeV CRs Cao et al. (2024b). Microquasars are x-ray binaries with relativistic jets Mirabel and Rodriguez (1994) and display phenomenology that is similar to active galactic nuclei (AGNs), although the size of the system is orders of magnitude different. It is suggested that shear and one-shot acceleration in the large-scale jet of radio galaxies can be responsible for the origin of UHECRs Kimura et al. (2018a); Caprioli (2015); Mbarek and Caprioli (2019); Mbarek et al. (2023, 2025). This motivates us to investigate whether the shear acceleration by microquasar jets can be Galactic PeVatrons. Previous studies also suggest microquasars as PeVatrons Cooper et al. (2020); Kimura et al. (2021); Yue et al. (2024), but these studies consider CR acceleration by shocks in jets or magnetic reconnections in accretion flows.
@@
The paper is organized as follows. In Sec. II, we study the physics of Galactic microquasars and the shear acceleration process in the jet-cocoon system of microquasars. In Sec. III, we propagate the escaped CRs from microquasars and compare the results with observations. In Sec. IV, we combine CRs from Galactic SNRs and contributions from the shear acceleration process from extragalactic clusters of galaxies and radio galaxies, and we proposed an overall picture of the origin of CRs on a wide energy range from GeV to $\sim 100\rm~EeV$. We give a summary and discussion in Sec. V.
@@ II.1 The physics of microquasar jets
Recently, Ref. Cao et al. (2024b) reported detection of gamma rays above 100 TeV from five microquasars, three of which (SS 433, V4641 Sgr, and GRS 1915+105) exhibited extended gamma-ray morphology whose emission zone is as large as $\sim 50-100$ pc . These gamma- ays are most likely originating from the radiation of nonthermal particles accelerated in the extended jets Alfaro et al. (2017); Aharonian et al. (2024); Cao et al. (2024b). The other two (Cyg X-1 and MAXI J1820+070) are consistent with point source, possibly suggesting a compact emission region Fang et al. (2024); Kuze et al. (2025), although the best fit positions of gamma-ray sources have slight offsets from those of x-ray binaries.
[…]
V4641 Sgr is another interesting microquasar that emits gamma-rays above 100 TeV. The morphology of the gamma-ray signal is clear two-sided lobe structure, and the size of the lobe is $\sim 100$ pc Cao et al. (2024b); Alfaro et al. (2024b). A spatially extended x-ray signature is also detected from this object, although the x-ray signal is concentrated on the central region within $\sim 15$ pc from the central black hole (Suzuki et al., 2025). The multiwavelength modeling of this system indicates $B\gtrsim 10\rm\mu G$ at the x-ray emission region.
GRS 1915+105 is a highly variable object that exhibits superluminal motion with extended jets of sub-pc scales Fender et al. (1999). Its Eddington ratio used to be $\sim 0.1-1$ (Done et al., 2004) and suddenly decreased to 0.01 in 2018 Koljonen and Hovatta (2021). The gamma-ray source detected by LHAASO exhibits an extension of $\sim 50$ pc Cao et al. (2024b), which is consistent with the two-sided lobe structure detected by ALMA Tetarenko et al. (2018).
The microquasar S26 in the Sculptor Galaxy NGC 7793 consists of a black hole and a Wolf-Rayet star. Its nebula spans $350\times 185\rm~pc$ Dopita et al. (2012), and the jet’s kinetic luminosity $L_{j}\sim 5\times 10^{40}\rm~erg~s^{-1}$ suggests super-Eddington accretion, making it a potential PeVatron  Abaroa et al. (2024).
@@
$\times{\left(\frac{L_{j}}{{10}^{39}~{\rm erg/s}}\right)}^{1/2}\left(\frac{\Gamma}{1.05}\right)^{-1}\left(\frac{\beta}{0.3}\right)^{1/2}$ (1)
where $Z$ is the nuclear charge, $\epsilon_{B}$ is the energy fraction carried by the magnetic field, $L_{j}$ is the total isotropic-equivalent luminosity of the jet, $\Gamma$ is the Lorentz factor, and $\beta$ is the characteristic velocity normalized by the speed of light. In our Galaxy, there are likely a few dozen black hole (BH) X-ray binaries with an x-ray luminosity of $L_{X}\gtrsim 10^{38}\rm~erg~s^{-1}$, each exhibiting sufficient luminosity, $L_{j}\sim 10^{39}\rm~erg~s^{-1}$, to potentially accelerate CRs up to the knee energy range. Here, we use frequently used assumptions of $L_{X}\sim 0.1\dot{M}c^{2}$ and $L_{j}\sim\dot{M}c^{2}$ to convert $L_{X}$ to $L_{j}$, where $\dot{M}$ is the mass accretion rate onto the BH.
@@ II.2 Shear acceleration in microquasar jets
Shear acceleration, a class of the Fermi acceleration, has been suggested to be an efficient mechanism to accelerate CRs, where charged particles gain energy inside an ordered shear velocity field Berezhko and Krymskii (1981); Earl et al. (1988); Ostrowski (1990); Webb (1989); Rieger and Duffy (2004); Rieger (2019); Matthews and Taylor (2021); Wan et al. (2025). It has been shown that the discrete shear acceleration Ostrowski (1990) occurring on the surface of the jet-cocoon system in radio galaxies may accelerate low-energy “seed” galactic CRs to the UHE energy range, and the predicted spectrum and composition are consistent with observations Kimura et al. (2018a); Seo et al. (2024).
In AGNs, the reacceleration may occur through the “one-shot” acceleration, where the seed CRs receive a boost of a factor of $\Gamma^{2}$ Gallant and Achterberg (1999), and it is possible for powerful radio galaxies with a Lorentz factor of $\Gamma\gtrsim 30$ to accelerate seed galactic CRs with energy $E\lesssim 10^{17}\rm~eV$ to the UHE energy range Caprioli (2015); Mbarek and Caprioli (2019). However, the one-shot acceleration is challenging in microquasars because of their low Lorentz factor jets, typically $\Gamma\lesssim 3$.
We consider the shear acceleration by cylindrical jets and cocoon as discussed in Kimura, Murase and Zhang (hereafter KMZ18) Kimura et al. (2018a). Based on our discussion in Sec. II.1, we assume that the radius of the jet is $r_{\rm jet}=3\rm~pc$, the length is $l_{\rm jet}=100\rm~pc$, the jet velocity normalized by the speed of light is $\beta_{j}=0.3$ (corresponding to Lorentz factor $\Gamma_{j}\simeq 1.05$), and the radius of the cocoon is $r_{\rm coc}=60\rm~pc$. The coherence length inside the cocoon is set to $l_{\rm coh}\sim 0.03r_{\rm coc}\sim 1.8\rm~pc$. The mean free path of CRs is given by $\lambda\sim l_{\rm coh}(E/E_{\rm coh})^{\delta}$, where $\delta=1/3$ for Kolmogorov turbulence and $E_{\rm coh}=ZeB_{\rm coc}l_{\rm coh}\sim 17Z(B_{\rm coc}/10{\rm~\mu G})(l_{\rm coh}/1.8\rm~pc)\rm~PeV$. CRs will undergo a discrete shear acceleration when the CR mean free path is longer than the width of the sheath layer, $\lambda\gtrsim r_{\rm sl}$. For $r_{\rm sl}\sim f_{\rm sl}r_{\rm jet}\sim 0.06(f_{\rm sl}/0.02)\rm~pc$, the injection energy is given by $\lambda\sim r_{\rm sl}$, which can be rewritten as
@@
$\sim$ $0.6\left(\frac{B_{\rm coc}}{\rm 10~\mu G}\right)\left(\frac{r_{\rm sl}}{0.06\rm~pc}\right)^{3}\left(\frac{l_{\rm coh}}{1.8\rm~pc}\right)^{-2}\rm~TeV.$
Thus, CRs with energies above several hundreds of GeV can undergo a discrete shear acceleration mechanism. In this context, injection energy is defined as the minimum kinetic energy required for a CR particle to undergo discrete shear acceleration. We should note that the width of the shear layer and injection energy to the discrete shear acceleration are still uncertain, and $f_{\rm sl}\sim 1$ is suggested by recent high-resolution MHD simulations when the kink instability is fully developed Wang et al. (2024).
Based on KMZ18 Kimura et al. (2018a), by balancing the acceleration and effective confinement timescales, the peak energy in the energy spectrum of escaping CR protons can be estimated to be
@@
$\simeq$ $4.3{\rm~PeV}\left(\frac{B_{\rm coc}}{10\mu\rm~G}\right)\left(\frac{l_{\rm coh}}{1.8\rm~pc}\right)^{1/2}\left(\frac{r_{\rm jet}}{3\rm~pc}\right)^{1/2}\left(\frac{\beta_{\rm jet}}{0.3}\right)\zeta,$
@@
$\times\left(\frac{r_{\rm coc}}{60\rm~pc}\right)^{-1}\left(\frac{v_{\rm exp}}{3000\rm~km~s^{-1}}\right),$
@@
$\times\left(\frac{r_{\rm coc}}{60\rm~pc}\right)^{-1}\left(\frac{B_{\rm coc}}{10\mu\rm~G}\right)\left(\frac{l_{\rm coh}}{1.8\rm~pc}\right)^{1/2}\left(\frac{\beta_{\rm jet}}{0.3}\right)\zeta\rm~erg~s^{-1}.$
Table 1: Model parameters used in Monte Carlo simulations.
@@
We perform Monte Carlo simulations with the code used by KMZ18 Kimura et al. (2018a) with parameters given in Table 1, and the resulting energy spectra and luminosity of the escaping CR protons are shown in Fig. 2. As seen in the figure, our shear acceleration scenario can achieve spectral peak around PeV energies with reasonable parameters (see Sec. V for a discussion about parameter choices). The escaping CR proton spectrum shows a hard component (spectral index $\sim$ 0.8) below the peak energy and a gradual cutoff above it.
Figure 2: The energy spectrum of escaping CRs predicted in the shear acceleration mechanism from Monte Carlo simulations. The parameters are listed in Table 1. The above plot is multiplied with $\rm exp(-\mathcal{R}/\mathcal{R}_{\rm max})$, where $\mathcal{R}_{\rm max}=4\times 10^{7}\rm~GeV$ is the maximum acceleration energy from the microquasar jets.